zarkam21:

For questions 11 - 14, solve the given triangles by finding the missing angle and other side lengths.

zarkam21:

1 attachment
zarkam21:

1 attachment
zarkam21:

1 attachment
Vocaloid:

sorry if this takes a while, people on QC are being butts as usual

Vocaloid:

anyway, for 11 you can use law of sines to solve for side B

zarkam21:

b/sin(B)=c/sin(C) b/sin(75)=7.8/

zarkam21:

ugh im confused

Vocaloid:

you don't have angle C or side c so we can't use those in the formula but we can use a and b, A and B, to find b b/sin(B) = a/sin(A) then you can solve for b

zarkam21:

b/sin(B) = a/sin(A) b/sin(75)=7.8/sin(48) b=10.14

Vocaloid:

cool now, since we have angle A and B, we can simply use the 180 - angle A - angle B rule to find angle C 180 - 75 - 48 = 57 for angle C then use law of sines again to find c

zarkam21:

c^2=a^2+b^2-2*a*b cos(C) c^2=7.8^2+10.14^2-2*7.8*10.14 (cos57) c^2=77.51 c=8.80

Vocaloid:

that's law of cosines

zarkam21:

oh shoot

zarkam21:

lol hold on im sorry

Vocaloid:

I mean, it'll probably work either way, but just to double check c/ (cos57) = 7.8/cos(48)

zarkam21:

6.35

Vocaloid:

huh that's weird anyway I'm going to check to see which one we should use, law of sines vs cosines

zarkam21:

okay sounds good il be here

Vocaloid:

apparently in terms of accuracy, law of cosines takes precedence over law of sines so I guess let's go with 8.80 for side c

Vocaloid:

anyway for 12) we only have one side so we have to start with law of sines first let's start by finding angle B = 180 - angle A - angle C = 122 then let's try finding side (a) b/sin(B) = a/sin(A)

zarkam21:

200/sin(21)=a/sin(37)

Vocaloid:

you used the side 200 which is across from angle B angle B is 122 not 21 so 200/sin(122) = a/sin(37) then find a

zarkam21:

141.93 =a

Vocaloid:

awesome then you can use law of cosines to find side c

zarkam21:

c²=a²+b²-2abCos(C) c^2=141.93^2+200^2-2*141.93*200 (cosC)

Vocaloid:

* 21 degrees

Vocaloid:

good but angle C = 21 degrees, plug that in and find c

zarkam21:

c²=a²+b²-2abCos(C) c^2=141.93^2+200^2-2*141.93*200 (cos21) c^2=7142.9 c=84.52

Vocaloid:

good, that's the triangle solved

Vocaloid:

anway for #13 let's label angle C and angle A on the triangle:

Vocaloid:



Vocaloid:

first let's find angle B by subtracting 180 - angle C - angle A = 180 - 18 - 140 = 22 degrees

zarkam21:

oh okay so now we have all the angles

zarkam21:

A=140 B=22 C=18

Vocaloid:

(got momentarily kicked from the site) anyway, since we only have 1 side, we have to use law of sines first to find side b b/sin(B) = a/sin(A) as usual, find b

Vocaloid:

* take a look at the diagram: angle A = 18 angle B = 22 angle C = 140

zarkam21:

b/sin(B) = a/sin(A) b/sin(22)=1/sin(128)

zarkam21:

b=0.48

Vocaloid:

b/sin(22)=1/sin(18) 18 not 128

zarkam21:

b=1.21

Vocaloid:

awesome now we got a, b, and all the angles, then use law of cosines to find c

zarkam21:

c²=a²+b²-2abCos(C) c^2=1^2+1.21^2-2*1*1.21 (cos140) c^2=4.32 c=2.08

Vocaloid:

awesome so that's triangle 13solved

zarkam21:

1 attachment
zarkam21:

so im trying to do #14 on my own. Im using #12 as a reference. But how did you get 122 in the first line

Vocaloid:

angle B = 180 - 21 - 37 = 122

zarkam21:

14/sin(28) = a/sin(37) a=17.95

zarkam21:

@Vocaloid i tagged you, so whenever you are back on we can continue

Vocaloid:

oh I didn't realize that there was a #14 whoops anyway the left side is good 14/sin(28) but side a is across from 112 so it would be 14/sin(28) = a/sin(112)

zarkam21:

14/sin(28) = a/sin(37) a=17.95 14/sin(28) 14/sin(28) = a/sin(112) a=27.65

zarkam21:

right

Vocaloid:

good, then you'd use law of cosines to find angle C

Vocaloid:

*side c

zarkam21:

what is b again

Vocaloid:

the side across from angle B

zarkam21:

14

zarkam21:

?*

Vocaloid:

yes

zarkam21:

c^2=a^2+b^2-2ab cosC c^2=27.65^2+14^2-2*27.65+14 (cos40) c^2=915.95 c=30.26

Vocaloid:

that plus sign between the 27.65 and the 14cos(40) shouldn't be there

zarkam21:

c^2=a^2+b^2-2ab cosC c^2=27.65^2+14^2-2*27.65*14 (cos40) c^2=367.45 c=19.17

Vocaloid:

yup that's it

zarkam21:

thankssss